# MMReview: A Multidisciplinary and Multimodal Benchmark for LLM-Based Peer Review Automation

### v2

@@ Introduction
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Open-source LLMs & MLLMs
@@
Gemini-2.5-flash-T -3.06 3.76 -3.58 3.86 -3.86 2.68 0.94 0.88 1.24 4.60 4.28 74.06 70.83 3.73 1.37 0.59
GLM-z1-flash -3.12 3.91 -3.70 3.84 -3.87 2.87 0.47 0.56 2.53 3.77 3.59 70.83 66.67 3.50 2.50 0.18
Table 2: Results on MMReview with text-only inputs.
@@
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Open-source MLLMs
Closed-source MLLMs
Gemini-2.5-flash-T -3.08 3.73 -3.56 3.88 -3.87 2.56 0.91 0.88 1.34 4.60 4.19 76.25 68.75 2.89 0.78 0.41
@@
Table 3: Results on MMReview with multimodal inputs.
Step Outcome Preference Attack
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$
Open-source MLLMs
Closed-source MLLMs
Gemini-2.5-flash-T -3.31 3.30 -3.67 3.74 -3.93 2.49 0.91 0.91 1.11 4.60 4.19 73.33 69.79 1.67 0.75
Table 4: Results on MMReview with pdf-as-img inputs.
Experiments
Evaluation details
We conduct extensive experiments on MMReview using a wide range of popular LLMs and MLLMs. Specifically, for open-source models, we include a diverse set of architectures and sizes, covering Qwen2.5VL (Bai et al. 2025), Kimi-VL-A3B (Team et al. 2025b), Qwen3 (Yang et al. 2025), InternVL3 (Zhu et al. 2025), OVIS2 (Lu et al. 2024), GLM-4.1V (Team et al. 2025a), and Deepseek (DeepSeek-AI et al. 2025). For advanced proprietary models, our selection includes GPT-4o, Gemini-2.5-Flash, GLM-Z1-Flash, and Claude-4-Sonnet.
As shown in Table 1, for tasks without objective evaluation metrics, namely S, SE, and WE, we employ BARTScore (Yuan, Neubig, and Liu 2021) and the ‘LLM-as-a-Judge’ paradigm (Bai et al. 2023; Zheng et al. 2023; Gu et al. 2024) to assess the similarity between model-generated and human-written review comments. For tasks where the model output is a numerical score, including SS, PS, CD, DD, CoD, FS, FW, and PI, we compute the Mean Absolute Error (MAE) between the model’s predicted score and the ground-truth score to quantify deviation. For classification-based tasks such as MD and PR, we evaluate performance using accuracy.
Main Results
The experimental results on MMReview, as shown in Tables 2, 3, and 4 (where T denotes Thinking and I denotes Instruct), yield the following key findings:
(1) Proprietary models consistently outperform open-source models. Closed-source models such as ChatGPT-4o, Claude, and Gemini demonstrate significant superiority across most metrics, particularly on high-level tasks such as Meta Decision (MD), Pairwise Ranking (PR), Conditional Decision (CD), and Chain-of-Thought Decision (CoD). This indicates that proprietary models exhibit greater consistency and reliability in generating review content, along with enhanced capabilities in understanding complex academic material and producing structured assessments.
(2) Input modality has a substantial impact on model performance. For most models, performance improves with the addition of visual inputs. Specifically, tasks directly related to evaluation outcomes (e.g., MD, PR, CD, DD, CoD) show notable gains under multimodal (text+image) input conditions. For instance, Qwen-2.5-VL-7B’s accuracy on the MD task increases from 72.92% to 75.83% with image input, its MAE on CD decreases from 2.43 to 2.20, and its PR accuracy rises from 59.38% to 64.58%. However, the PDF-as-image modality poses challenges: most models—including advanced proprietary ones like Claude and Gemini—perform worse on this modality than on structured text-image input, though closed-source models remain comparatively more robust than open-source counterparts.
(3) Proprietary models exhibit stronger robustness to prompt injection. On the Prompt Injection (PI) task, proprietary models achieve MAE values below 0.6 in most cases, often under 0.5, indicating consistent judgment with and without injected prompts. In contrast, open-source models exhibit highly variable performance under prompt injection, suggesting greater susceptibility to adversarial manipulation.
(4) Larger model size does not guarantee superior performance. A comparative analysis of the Qwen2.5-VL and OVIS2 model families reveals that performance does not uniformly improve with scale. In several cases, smaller models outperform their larger counterparts, suggesting that model size is not a definitive determinant of effectiveness in peer review tasks.
More analysis of the result can be found in the supplementary materials.
Further Analysis
Are Thinking Models More Suitable for Review?
For closed-source models, performance differences between Gemini-2.5-Flash and Gemini-2.5-Flash-Thinking are minimal. In contrast, within the open-source Kimi-VL-A3B and Deepseek series, a more nuanced trend emerges: while non-thinking variants may outperform on some step-based tasks, thinking models consistently excel in outcome-based and preference-based tasks critical to review decisions. This indicates that thinking-enhanced models are better suited for generating final peer review judgments. Further supporting evidence is provided in the supplementary material.
@@
More further analysis can be found in the supplementary materials of our submission.
Conclusion
In this work, we present MMReview—a multidisciplinary and multimodal benchmark designed to evaluate the capabilities of Large Language Models (LLMs) in academic peer review. The benchmark encompasses 4 thematic categories and 13 distinct tasks. Its core features include coverage across diverse academic disciplines, support for multimodal input formats, and a comprehensive set of evaluation tasks that span the full peer review pipeline. Leveraging MMReview, we conducted extensive evaluations of state-of-the-art LLMs and MLLMs. We envision MMReview as a standardized evaluation platform that can catalyze the development of more efficient LLM-assisted peer review systems. In future work, we plan to incorporate larger-scale datasets specifically curated for training LLMs in the context of scholarly review.
Appendix
As a supplement, we provide additional materials in the appendix, including more detailed statistics on the benchmark and tasks (Appendix A), extended experimental results and analyses (Appendix B), implementation details of the experiments (Appendix C), several illustrative case studies (Appendix D), as well as the prompts used during the construction and application of the benchmark (Appendix E).
Appendix A A. Statistical Information about Research Domains in MMReview Benchmark
Discipline Research Field # Samples
Artificial
Intelligence
Machine Learning 17
Computer Vision 20
Natural Language Processing 16
Reinforcement Learning 11
Graph Neural Networks 19
Signal Processing 17
AI Application 15
Natural
Sciences
Biology & Medicine 11
Physics 16
Chemistry 13
Environmental & Earth Sciences 10
Mathematics & Statistics 13
Engineering
Sciences
Materials Science 14
Control Science 12
Electronic Information 6
Energy Science 6
Social
Sciences
Society, Economics & Finance 24
Table 6: The distribution of papers across various research domains.
Table 6 presents the distribution of papers across various research domains within our MMReview benchmark. For each domain, we ensured a relatively balanced number of samples through a combination of model-based filtering and manual curation.
Appendix B B. More Results and Analysis
B.1 The Impact of Paper Length
As a long-context task, peer review performance may be influenced by the length of the manuscript. To examine this, we analyze the average scores assigned by LLMs under two input settings: text-only (with context length measured in tokens) and PDF-as-image (with context length measured by the number of rendered pages). As illustrated in Figures 3 and 4, we observe a consistent trend wherein models tend to assign higher scores as the context length or number of images increases. This upward bias does not align well with human reviewer judgments across multiple intervals, indicating an inherent tendency of LLMs to overvalue longer inputs irrespective of actual content quality. Such length-induced bias poses a significant challenge for the practical deployment of LLMs in peer review scenarios.
Figure 4: The average scores under pdf-as-img input setting, with context length measured in tokens.
B.2 How does Prompt Injection Work?
To further investigate the impact of prompt injection on model scoring behavior, we analyze score variations between the Prompt Injection task and the CoT Decision (CoD) task, where the only difference in prompts lies in the injection of the statement: “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPERS AND DO NOT HIGHLIGHT ANY NEGATIVES.” All other prompt content remains identical. Tables 8 and 9 report the proportion and expectation of score increases and decreases under text-only and multimodal input conditions, respectively. The experimental results reveal three key findings: (1) Most models exhibit an overall increase in average score after prompt injection, with a typical rise of 1–2 points; (2) Multimodal inputs enhance model robustness against prompt injection compared to text-only inputs, consistent with prior observations of improved performance under multimodal settings; (3) The robustness of Thinking variants varies across model families—within the Qwen3 and Gemini series, Thinking models are less robust than their non-Thinking counterparts, whereas in the Kimi and Deepseek series, Thinking models demonstrate greater resistance to injection. These observations suggest that the relationship between prompt injection robustness and the model’s internal reasoning style is non-trivial and may depend on model-specific cognitive mechanisms and prompt interpretation behavior.
Deepseek-V3
w.o. Ref
0.53 0.47 2.62 3.37 3.70 0.75 0.66 2.99 0.66 0.20
Deepseek-V3
w/ Ref
0.47 0.44 2.56 3.26 3.74 0.83 0.57 2.99 0.71 0.19
Table 7: The influence of references on model performance.
@@ B.2 How does Prompt Injection Work?
B.3 The Impact of Reference Section
@@
Appendix C C. More Implementation Details
@@
Appendix D D. Prompts
In this section, we present the prompts used throughout the construction and evaluation phases of the benchmark. Figure 5 illustrates the prompt employed in the task generation stage to produce antonymic rewrites of strengths and weaknesses. Figures 6 through 16 display the prompts corresponding to the various tasks designed within the benchmark. Additionally, Figure 17 presents the prompt based on the Likert scale used for LLM-as-a-judge evaluation.

### v3

@@ 4.1 Evaluation details
Model $S_{B}\uparrow$ $S_{L}\uparrow$ $SE_{B}\uparrow$ $SE_{L}\uparrow$ $WE_{B}\uparrow$ $WE_{L}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
InternVL3-2B -3.03 3.64 -3.66 3.56 -3.95 2.15 0.47 0.47 2.35 3.43 3.29 66.25 53.13 3.14 1.37 1.11
Qwen2.5-VL-3B -3.23 3.58 -3.89 3.36 -4.00 1.67 0.00 0.00 1.46 4.37 4.46 64.91 73.96 3.46 2.95 0.05
@@
[…]
Table 3: Results on MMReview with text-only inputs, where T denotes Thinking and I denotes Instruct.
Model
Size
@@
$S_{B}\uparrow$ $S_{L}\uparrow$ $SE_{B}\uparrow$ $SE_{L}\uparrow$ $WE_{B}\uparrow$ $WE_{L}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Multimodal
Tiny -3.12 3.59 -3.76 3.37 -3.97 2.00 0.48 0.46 2.14 3.24 3.56 59.66 56.77 2.95 1.11 0.42
Small -3.16 3.50 -3.72 3.49 -3.99 2.13 0.48 0.46 2.41 3.64 3.82 69.72 61.98 3.00 1.77 0.31
Middle -3.05 3.71 -3.67 3.63 -3.94 2.35 0.72 0.69 2.21 3.66 3.86 73.20 61.21 3.00 1.55 0.47
Large -3.07 3.82 -3.59 3.86 -3.86 2.78 0.76 0.75 1.59 3.48 3.34 78.79 65.97 3.25 1.37 0.40
PDF-as-img
Tiny -3.25 3.27 -4.00 2.98 -4.11 1.83 0.47 0.73 1.86 2.93 3.59 67.47 54.95 2.35 1.08 \
Small -3.55 2.89 -3.88 3.20 -4.08 1.99 0.47 0.44 2.14 3.38 3.47 72.78 59.17 2.99 1.72 \
Middle -3.14 3.55 -3.70 3.54 -3.95 2.29 0.68 0.70 2.06 3.65 3.76 71.77 55.90 3.09 1.34 \
Large -3.28 3.50 -3.67 3.72 -3.90 2.68 0.77 0.75 1.39 3.53 3.36 77.99 63.89 2.85 1.60 \
@@
Table 4: Results on MMReview with multimodal and pdf-as-image inputs. The detailed results from models can be found in Table 8 and 9 in Appendix A.
4.2 Main Results
We present the performance metrics for each model in the text-only input mode in Table 3. To observe the performance across different model categories, we calculated the average performance for each group. Table 4 displays the average performance metrics for each model group under the multimodal and PDF-as-image input mode. The specific performance of each model is provided in Tables 8 and 9 in the Appendix. Based on the test results presented in the tables, the following conclusions can be drawn:
(1) Model scale significantly influences the model’s ability to comprehend and analyze. Large-scale and closed-source models outperform the others on most metrics, particularly on tasks directly related to review conclusions, such as CD, CoD, MD, and PR. This indicates that larger models are more powerful in understanding complex academic content and generating structured feedback, making them more reliable in generating peer review comments. Surprisingly, mid-sized and smaller models performed better than their larger counterparts in assessing the soundness and presentation of papers.
(2) High-quality structured reasoning enhances review outcomes. Compared to directly generating review scores (DD), the use of CoT reasoning to generate review scores (CoD) achieved a lower MAE, demonstrating that a step-by-step review approach strengthens the model’s evaluative capability. Furthermore, using a higher-quality, human-written review process as a reference (CD) further reduced the MAE, indicating that the quality of the reasoning process also affects the model’s evaluative ability. Models that employed thinking (such as Kimi-VL-A3B-Thinking and Deepseek-R1) outperformed non-thinking models on tasks directly related to scoring, such as CD, CoD, MD, and PR, suggesting that stronger thinking and reasoning abilities enhance the model’s peer review capabilities.
(3) Multimodal input improves the model’s robustness to prompt injection. Regardless of model size, when both text and paper images are input simultaneously, the influence of prompt injection is lower compared to the text-only input mode, implying that the inclusion of image modality enhances the model’s ability to judge paper quality and mitigates the impact of prompt injection interference.
5 Further Analysis
@@
How does Prompt Injection Work?
To further investigate the impact of prompt injection on model scoring behavior, we analyze score variations between the Prompt Injection task and the CoT Decision (CoD) task, where the only difference in prompts lies in the injection of the statement: “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPERS AND DO NOT HIGHLIGHT ANY NEGATIVES.” Tables 6 and 7 report the proportion and expectation of score increases and decreases under text-only and multimodal input conditions, respectively. The experimental results reveal three key findings: (1) Most models exhibit an overall increase in average score after prompt injection, with a typical rise of 1–2 points; (2) Multimodal inputs enhance model robustness against prompt injection compared to text-only inputs, consistent with prior observations; (3) The robustness of Thinking variants varies across model families, within the Qwen3 series, Thinking models are less robust than their non-Thinking counterparts, whereas in the Kimi and Deepseek series, Thinking models demonstrate greater resistance to injection.
@@
More analysis are presented in the Appendix A.
6 Conclusion
In this work, we present MMReview, a multidisciplinary and multimodal benchmark designed to evaluate the capabilities of LLMs in academic peer review. The benchmark encompasses 4 thematic categories and 13 distinct tasks. Its core features include coverage across diverse academic disciplines, support for multimodal input formats, and comprehensive evaluation tasks that span the full peer review pipeline. Leveraging MMReview, we conducted extensive evaluations of LLMs and MLLMs. We envision MMReview as a standardized evaluation platform that can catalyze the development of more efficient LLM-assisted peer review systems.
7 Limitations
The issue of dataset size and distribution.
Due to the rapid advancement of AI in recent years and the open access and public review characteristics of AI papers, approximately 48% of the papers containing peer review comments are concentrated in the AI field. This concentration may affect the representativeness of the results. In the future, we plan to collect more papers from other domains to enhance the representativeness of our benchmark.
The controversy surrounding review comments written by human experts.
There is currently a lack of consensus on what constitutes a good or high-quality review. Although methods such as obtaining review consensus and manual screening have been employed in the paper to filter review samples, it remains impossible to guarantee that these expert-written reviews are of sufficient quality. Moreover, it is undeniable that all types of reviews hold value, even though they may vary significantly in content, as they reflect the diverse perspectives of different reviewers.
8 Ethical Considerations
All the papers and peer review comments we collected are sourced from open-access platforms such as OpenReview, NeurIPS, and Nature. These platforms state that the content they publish, including but not limited to the papers themselves and their peer review comments, is licensed or permitted for research purposes under the Creative Commons Attribution International 4.0 license. We ensure that the collection and processing of these papers and reviews are conducted for research purposes and comply with the copyright agreements of the platforms.
Our research on the ability of LLMs to generate peer review comments does not advocate for the complete replacement of human reviewers with LLMs, as this might open the door to potential misuse and manipulation. Instead, we envision that in the current era of a proliferation of academic papers, LLMs can serve as an auxiliary tool. Similar to the practices already adopted by academic conferences like ICLR and AAAI, peer review comments generated by LLMs could be used as references to help reduce the workload of human reviewers to some extent.
Appendix
As a supplement, we provide additional materials in the appendix, including extended experimental results and analyses (Appendix A), implementation details of the experiments (Appendix B), case studies C, as well as the prompts used during the construction and application of the benchmark (Appendix D).
Step Outcome Preference Attack
Model $S_{B}\uparrow$ $S_{L}\uparrow$ $SE_{B}\uparrow$ $SE_{L}\uparrow$ $WE_{B}\uparrow$ $WE_{L}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
InternVL3-2B -3.11 3.67 -3.95 2.87 -4.04 1.89 0.52 0.48 1.99 2.56 2.94 60.94 53.13 2.48 1.10 0.57
Qwen2.5-VL-3B -3.11 3.55 -3.74 3.33 -3.97 1.89 0.47 0.44 1.87 3.24 3.74 55.42 57.29 3.18 2.88 0.08
[…]
Table 8: Results on MMReview with multimodal inputs.
Step Outcome Preference Attack
Model $S_{B}\uparrow$ $S_{L}\uparrow$ $SE_{B}\uparrow$ $SE_{L}\uparrow$ $WE_{B}\uparrow$ $WE_{L}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$
InternVL3-2B -3.28 3.28 -4.80 1.51 -4.55 1.19 0.48 1.61 1.57 1.77 3.34 72.38 50.00 3.00 1.11
Qwen2.5-VL-3B -3.29 3.12 -3.76 3.34 -3.98 1.86 0.47 0.44 1.73 3.05 3.72 53.75 54.17 3.01 2.94
[…]
Table 9: Results on MMReview with pdf-as-image inputs.
Appendix A More Results and Analysis
A.1 The Impact of Paper Length
As a long-context task, peer review performance may be influenced by the length of the manuscript. To examine this, we analyze the average scores assigned by LLMs under two input settings: text-only (with context length measured in tokens) and PDF-as-image (with context length measured by the number of rendered pages). As illustrated in Figures 2 and 3, we observe a consistent trend wherein models tend to assign higher scores as the context length or number of images increases. This upward bias does not align well with human reviewer judgments across multiple intervals, indicating an inherent tendency of LLMs to overvalue longer inputs irrespective of actual content quality. Such length-induced bias poses a significant challenge for the practical deployment of LLMs in peer review scenarios.
Figure 3: The average scores under pdf-as-image input setting, with context length measured in the number of images.
Deepseek-V3
w.o. Ref
0.53 0.47 2.62 3.37 3.70 0.75 0.66 2.99 0.66 0.20
Deepseek-V3
w/ Ref
0.47 0.44 2.56 3.26 3.74 0.83 0.57 2.99 0.71 0.19
Table 10: The influence of references on model performance.
A.2 The Impact of Reference Section
@@
Appendix B More Implementation Details
@@
All of the human annotators and reviewers mentioned in this paper were selected from a group of five PhD students, each with extensive submission experience and a background in reviewing for academic conferences such as ARR and AAAI, as well as an adequate knowledge base in the fields covered by the papers. These individuals were provided with clear instructions regarding the high-quality paper selection and review comment annotation tasks they were required to complete, along with the objectives of these tasks. They were compensated at a market-average hourly rate of \$30/h for their work.
Appendix C Case studies
In this section, we present two representative case studies from the evaluation results of chatgpt-4o-latest, corresponding to papers submitted to ICLR and Nature Communications, respectively.
In Case 1 (Figures 49), the model was tasked with completing all benchmark tasks using a text-only input. Human annotations highlight (in green) the portions of the model’s responses that align with the original reviewer comments. The evaluated paper was a rejected submission to ICLR 2024. While the model provided a relatively favorable overall assessment and high score in the CoT Decision task, it identified more weaknesses than strengths, consistent with the human reviewers’ concerns, indicating partial alignment with human judgment. When acting as an area chair in the Meta Decision task, the model successfully synthesized reviewer opinions to arrive at a justified rejection decision, showcasing its ability to integrate and summarize multiple reviews. However, in the Prompt Injection (PI) task, despite the prompt being identical to that of the CoD task, the model’s output exhibited a strong bias toward highlighting strengths (highlighted in yellow) and delivered a more favorable final assessment, underscoring the significant influence of prompt injection on model behavior.
In Case 2 (Figures 1013), the model reviewed an accepted paper from Nature Communications using the pdf-as-image input modality. Again, human annotations indicate alignment between the model’s outputs and human-written review comments. As expected for an accepted paper, the model emphasized strengths over weaknesses. However, in the Meta Decision task, due to the absence of explicit quantitative scores in the original reviewer comments, the model struggled to infer the reviewers’ overall preference and mistakenly issued a rejection decision. This case highlights the challenges LLMs face when interpreting unstructured or qualitative feedback in the absence of scoring signals.
Appendix D Prompts
In this section, we present the prompts used throughout the construction and evaluation phases of the benchmark. Figure 14 illustrates the prompt employed in the task generation stage to produce antonymic rewrites of strengths and weaknesses. Figures 15 through 25 display the prompts corresponding to the various tasks designed within the benchmark. Additionally, Figure 26 presents the prompt based on the Likert scale used for LLM-as-a-judge evaluation.
Figure 4: Result of Summary task in case 1.
Figure 5: Result of SE task in case 1.
Figure 6: Result of WE task in case 1.
Figure 7: Result of MD task in case 1.
Figure 8: Result of CoD task in case 1.
Figure 9: Result of PI task in case 1.
Figure 10: Result of Summary task in case 2.
Figure 11: Result of SE task in case 2.
Figure 12: Result of WE task in case 2.
Figure 13: Result of MD task in case 2.